OpenStudy (anonymous):

HELP PLEASE! I need help with part B and part C. The table shows the outputs y for different inputs x: Input (x) 2 4 6 8 Output (y) 1 2 3 4 Part A: Do the data in this table represent a function? Justify your answer. (3 points) Part B: Compare the data in the table with the relation f(x) = 3x – 10. Which relation has a greater value when x = 8? (2 points) Part C: Using the relation in Part B, what is the value of x if f(x) = 80? (5 points)

OpenStudy (jdoe0001):

so.. what did you get for part A?

OpenStudy (anonymous):

That it is a function.Its a function because for every domain theres a range.You can only have one range for each domain.

OpenStudy (anonymous):

divide input by 2 see what you get?

OpenStudy (jdoe0001):

right.... well, for part B is only asking, what's "y" when x = 8 from the table data? and what's "y" when x = 8 for the expresion f(x)=y = 3x-10? which one is biggers?

OpenStudy (jdoe0001):

\(\begin{array}{llll} x&y \\\hline\\ 2&1\\ 4&2\\ 6&3\\ {\color{red}{ 8}}&{\color{blue}{ 4}} \end{array}\qquad \qquad \begin{array}{llll} f(x)={\color{blue}{ y}}=3{\color{red}{ x}}-10 \\ \quad \\ f(x)={\color{blue}{ y}}=3{\color{red}{ (8)}}-10 \end{array} \)

OpenStudy (jdoe0001):

so... notice what "y" is for the data table, is 4, when x =8 what about the 3x-10? when x = 8, what is "y"?

OpenStudy (anonymous):

y = f(x) = x/2

OpenStudy (anonymous):

what @jdoe0001 is struggling to say is that a function is a special type of relation in which there is a one-to one correspondence between the x(input) and the y(output)